# Anchoring and Mixed-Norm Contractions in Averaging-Learning Dynamics

### v1

@@ 1.2 Notation and preliminaries
where $1\leq p\leq\infty$. In particular, if $p=1$ or $p=\infty$, we have the following formulae of the induced matrix $1$-norm and $\infty$-norm:
@@
  • $M$ is row stochastic if each row-sum of $M$ is equal to $1$.
  • $M$ is row substochastic if each row-sum of $M$ is at most $1$.
  • $M$ is proper substochastic if $M$ is row substochastic and at least one row-sum of $M$ is less than $1$.
@@ 1.3 Directed graphs and row-stochastic matrices
In this paper, we require some basic facts about directed graphs. A directed graph (or digraph) $G$ is an ordered pair $(V,E)$ consisting of the vertex set $V$ together with the set of arcs $E$. Each arc is an ordered pair $(i,j)$ where $i,j\in V$. In our visual representation of digraphs, an arc $(i,j)$ is represented by an arrow starting from the initial vertex $i$ and ending at the terminal vertex $j$. The arc $(i,i)$ is called a self-loop at vertex $i$. The outdegree of a vertex $v\in V$ is the number of arcs $(i,j)\in E$ such that $v=i$. Likewise, the indegree of a vertex $v\in V$ is the number of arcs $(i,j)\in E$ such that $v=j$. Let $G=(V,E)$ be a digraph and let $W\subseteq V$. The digraph induced by $W$ is the digraph $G|_{W}=(W,E|_{W})$ where $E|_{W}$ is the set of arcs $(i,j)$ such that $(i,j)\in E|_{W}$ if and only if $i,j\in W$.
Let $k\geq 0$ be an integer. A walk $\omega$ starting from vertex $i_{0}$ and ending at vertex $i_{k}$ in a digraph $G=(V,E)$ is a sequence of vertices $\langle i_{0},i_{1},\dots,i_{k}\rangle$ such that $(i_{t},i_{t+1})\in E$ for all $t\in\{0,1,\dots,k-1\}$. The length of the walk $\omega=\langle i_{0},i_{1},\dots,i_{k}\rangle$, denoted by $\lvert\omega\rvert$, is equal to $k$. We can also write the walk $\omega=\langle i_{0},i_{1},\dots,i_{k}\rangle$ as $i_{0}\to i_{1}\to\dots\to i_{k}$. Note that a walk may include a self-loop, or a vertex can be visited more than once. A walk that starts and ends at the same vertex is called a closed walk. For example, we consider any walk of length 0 to be a closed walk. A cycle is a closed walk of positive length with no repeated vertex. We consider a self-loop as a cycle of length $1$. A directed graph with no cycle is called a directed acyclic graph (DAG).
Let $G=(V,E)$ be a digraph. We define a relation on $V$ where $i\sim j$ if and only if there is a walk in $G$ from $i$ to $j$, and there is also a walk from $j$ to $i$. This relation is an equivalence relation on $V$. An equivalence class induced by this equivalence relation is called a strongly connected component (SCC) of $G$. If $G$ consists of only one SCC, then we call $G$ strongly connected. Equivalently, $G$ is strongly connected if and only if for all $i,j\in V$, there exists a walk in $G$ from $i$ to $j$. Let $W$ be an SCC of $G$. Clearly, the induced digraph $G|_{W}$ is strongly connected. We call $W$ a sink SCC if for all $(i,j)\in E$, $i\in W\implies j\in W$. In words, $W$ is a sink if there is no arc with its initial vertex in $W$ and its terminal vertex outside of $W$. For example, any vertex that has outdegree zero forms a sink SCC.
@@
Let $A$ be a row-stochastic matrix. The behavior of $\lim_{t\to\infty}A^{t}$ can be broadly categorized into two types:
@@
The following theorem is a combinatorial characterization of the existence of $\lim_{t\to\infty}A^{t}$ based on the underlying digraph of a given row-stochastic $A$. Note that a row-stochastic matrix $A$ is condensely aperiodic if and only if for all sink SCC $W$ of $G[A]$, the induced digraph $G[A]\,|_{W}$ is aperiodic.
@@
Let $A$ be a row-stochastic matrix of positive size. Then $\lim_{t\to\infty}A^{t}=M$ where $M$ is row stochastic, if and only if $A$ is condensely aperiodic.
@@
Let $A$ be a row-substochastic matrix of positive size. If $A$ is condensely aperiodic then $\lim_{t\to\infty}A^{t}=M$ where $M$ is row substochastic.
@@
Suppose that $A$ is condensely aperiodic. Then for each $i$, $S_{i}=0$ or $S_{i}$ is aperiodic. Suppose $S_{i}$ is aperiodic and thus, primitive. Suppose also that $S_{i}$ is row stochastic, i.e., $\rho(S_{i})=1$. By the Perron–Frobenius theorem [48, Theorem 1.1], $S_{i}$ has $1$ as a simple eigenvalue, and for all eigenvalues $\lambda\neq 1$ of $S_{i}$, we have $\lvert\lambda\rvert<1$. The eigenvalues of $A$ are precisely the eigenvalues of $U_{1},U_{2},\dots,U_{k},S_{1},S_{2},\dots,S_{m}$. Let $\lambda$ be an eigenvalue of $A$. Obviously, $\lvert\lambda\rvert\leq 1$. Suppose that $\lvert\lambda\rvert=1$. Then at least one of $S_{1},S_{2},\dots,S_{m}$ has $\lambda$ as an eigenvalue. Suppose that for some $i$, $S_{i}$ has $\lambda$ as an eigenvalue. Since $\lvert\lambda\rvert=1$ then $\rho(S_{i})=1$. This implies that $S_{i}$ must be aperiodic and row stochastic. Hence, it is clear that $\lambda=1$. Furthermore, as an eigenvalue of $A$, $\lambda=1$ is semisimple: its algebraic and geometric multiplicities are equal. By [31, Theorem 2.1], we conclude that there exists a matrix $M$ such that $\lim_{t\to\infty}A^{t}=M$. Moreover, observe that all entries of the vector $A^{t}\mathbf{1}$ is at most $1$, which is denoted as $A^{t}\mathbf{1}\leq\mathbf{1}$. Taking limit yields $M\mathbf{1}\leq\mathbf{1}$ and thus, $M$ is also row substochastic. ∎
@@ 2 Time-invariant case and the spectral radius
In the time-invariant case, we have that $A_{t}=A$ and $\mathcal{E}_{t}=\mathcal{E}$ for all integers $t\geq 0$. We fix the notation $A$ and $\mathcal{E}$ for the rest of Section 2, unless stated otherwise. Next, Equation (2.1) becomes
@@
By Theorem 1.1, AA-$\mathcal{E}$ is zero-convergent if and only if $\rho(A-\mathcal{E})<1$. Even if $\rho(A-\mathcal{E})\geq 1$, it is possible that consensus can still be eventually reached, depending on $X_{0}-\bar{\sigma}$. An example is if $\lim_{t\to\infty}(A-\mathcal{E})^{t}=M$, where $M\neq 0$ is singular. Our main focus in this section is to derive various conditions for the entries of $A$ and $\mathcal{E}$ such that $\rho(A-\mathcal{E})<1$.
@@
We know that $\rho(A-\mathcal{E})\leq\lVert A$-$\mathcal{E}\rVert_{\infty}$. Suppose that $0<\mathcal{E}(i)<2A(i,i)$ for all $i$ in $\{1,2,\dots,n\}$. We claim that $\lVert A-\mathcal{E}\rVert_{\infty}<1$. Since $A$ is row stochastic, for each row $i\in\{1,2,\dots,n\}$, we have that
@@ 2.1 Anchor agents and the condensely-anchored property
Let $A$ be a row-stochastic matrix and let $\mathcal{E}$ be a diagonal nonnegative matrix of the same size $n>0$. Let $i\in\{1,2,\dots,n\}$. An agent $i$ is called an anchor of AA-$\mathcal{E}$ if $0<\mathcal{E}(i)<2A(i,i)$. Alternatively, the row/index/vertex $i$ is also called an anchor of AA-$\mathcal{E}$ if $0<\mathcal{E}(i)<2A(i,i)$. If $\mathcal{E}(i)=0$ then an agent $i$ is called defective. Consequently, an agent $i$ is a non-anchor if either $i$ is defective or $\mathcal{E}(i)\geq 2A(i,i)$.
@@
Let $A$ be a row-stochastic matrix of size $n>0$, and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$ such that $A-\mathcal{E}\geq 0$. Let $i\in\{1,2,\dots,n\}$. The following are equivalent:
@@
    The $i$-th row-sum of AA-$\mathcal{E}$ is strictly less than $1$.
@@
Let $B$ be a row-substochastic matrix of size $n>0$. A row/index $i\in\{1,2,\dots,n\}$ is an anchor of $B$ if and only if the $i$-th row-sum of $B$ is strictly less than $1$.
@@
Observe that there exists a unique pair of row-stochastic matrix $A$ and diagonal nonnegative matrix $\mathcal{E}$, both of size $n$, such that $B=A$-$\mathcal{E}$. The conclusion then immediately follows from $(1)\iff(3)$ in Proposition 2.4. ∎
By Corollary˜2.5, it is clear that any row-stochastic matrix $A$ does not have any anchor. Outside the context of this paper, Corollary˜2.5 can instead be used as the starting point to define anchors for row-substochastic matrices.
@@
Let $A$ be a row-stochastic matrix and let $\mathcal{E}$ be a diagonal nonnegative matrix of the same size. We call AA-$\mathcal{E}$ condensely anchored if every sink SCC of $G[A-\mathcal{E}]$ contains an anchor of AA-$\mathcal{E}$.
Note that Definition 2.6 above parallels that of condensely aperiodic. Clearly, if each row of AA-$\mathcal{E}$ is an anchor, then AA-$\mathcal{E}$ is condensely anchored. On the other hand, if AA-$\mathcal{E}$ does not have any anchor, then AA-$\mathcal{E}$ is not condensely anchored. For example, the row-stochastic matrix $A$, if written as A0A-0, is not condensely anchored. Hence, this emphasizes that the definition of condensely anchored applies to AA-$\mathcal{E}$ instead of $A$.
@@
Let $A$ be a row-stochastic matrix and let $\mathcal{E}$ be a diagonal nonnegative matrix of the same size $n>0$. Then AA-$\mathcal{E}$ is condensely anchored if and only if for all $i\in\{1,2,\dots,n\}$, there exists a walk in the digraph $G[A-\mathcal{E}]$ from vertex $i$ to some anchor vertex of AA-$\mathcal{E}$.
@@
Let $G=(V,E)$ be a digraph. Instead of involving anchors as defined in Definition˜2.3, we can simply fix any subset $U\subseteq V$. Hence, we can jointly reformulate both Definition 2.6 and Proposition 2.7 as: Every sink SCC of $G$ contains a vertex from $U$ if and only if for all $i\in V$, there exists a walk in $G$ from $i$ to some vertex in $U$. This statement is inherently graph-theoretic, as it can be stated independently of the matrices $A$ and $\mathcal{E}$.
@@ 2.2 The case where AA-$\mathcal{E}$ is nonnegative
Let $B$ be a row-substochastic matrix of size $n>0$. Let $i\in\{1,2,\dots,n\}$ and let $\widehat{W}_{i}(B)$ be the set of walks $\omega$ in $G[B]$ such that $\omega$ starts at $i$ and ends at some vertex $j$, where the $j$-th row-sum of $B$ is strictly less than $1$. We define the index of contraction of $B$ as
@@
be a row-substochastic matrix. By Corollary˜2.5, note that only row $4$ of $B$ is an anchor. There exists a unique pair of row-stochastic matrix
@@
Figure 7: The underlying digraph of both matrices $A$ and $B$ in Example˜2.10.
@@
Let $A$ be a row-stochastic matrix of size $n>0$, and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$ such that $A-\mathcal{E}\geq 0$. Then, the following are equivalent:
@@
$(1)\iff(2)$ follows directly from Proposition˜2.7. By [4, Corollary 2.6] or [60, Theorem A.1.6], we obtain $(3)\iff(4)$. We will now show that $(2)\iff(3)$. First, we know that AA-$\mathcal{E}$ is row substochastic. From Definition˜2.9, we have that $\widehat{W}_{i}(A-\mathcal{E})$ is the set of walks $\omega$ in $G[A-\mathcal{E}]$ such that $\omega$ starts at $i$ and ends at some vertex $j$, where the $j$-th row-sum of AA-$\mathcal{E}$ is strictly less than $1$. By Proposition 2.4, we also have that a vertex $j$ is an anchor of AA-$\mathcal{E}$ if and only if the $j$-th row-sum of AA-$\mathcal{E}$ is strictly less than $1$. Hence, we obtain
@@
Let $B$ be a row-substochastic matrix of size $n>0$. We know that there exists a unique pair of row-stochastic matrix $A$ and diagonal nonnegative matrix $\mathcal{E}$, both of size $n$, such that $B=A$-$\mathcal{E}$. Hence, Theorem˜2.11 can also be stated in terms of row-substochastic matrix $B$ instead of AA-$\mathcal{E}$.
@@
In Example 2.13, we fix a row-stochastic matrix $A$, and we consider two diagonal nonnegative matrices $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$. For each matrix $A-\mathcal{E}_{1}$ and $A-\mathcal{E}_{2}$, there are four defective agents but only a single anchor.
@@
The digraphs $G[A]$, $G[A-\mathcal{E}_{1}]$, and $G[A-\mathcal{E}_{2}]$ are all identical and condensely aperiodic, see Table˜1. Moreover, $\{1\}$ is the only sink SCC of each of these digraphs. Vertex $1$ of $G[A-\mathcal{E}_{1}]$ is its sole anchor, while vertex $5$ of $G[A-\mathcal{E}_{2}]$ is its sole anchor. In Table 1, we can see that $A-\mathcal{E}_{1}$ is condensely anchored, while $A-\mathcal{E}_{2}$ is not condensely anchored. By Theorem 2.11, we have $\lim_{t\to\infty}(A-\mathcal{E}_{1})^{t}=0$, while $A-\mathcal{E}_{2}$ is not zero-convergent. We find that

### v2

# arXiv:2602.22627v2[math.DS] 28 Feb 2026

@@ 1.2 Notation and preliminaries
where $1\leq p\leq\infty$. In particular, if $p=1$ or $p=\infty$, we have the following formulae of the induced matrix 11-norm and $\infty$-norm:
@@
  • $M$ is row stochastic if each row-sum of $M$ is equal to 11.
  • $M$ is row substochastic if each row-sum of $M$ is at most 11.
  • $M$ is proper substochastic if $M$ is row substochastic and at least one row-sum of $M$ is less than 11.
@@ 1.3 Directed graphs and row-stochastic matrices
In this paper, we require some basic facts about directed graphs. A directed graph (or digraph) $G$ is an ordered pair $(V,E)$ consisting of the vertex set $V$ together with the set of arcs $E$. Each arc is an ordered pair $(i,j)$ where $i,j\in V$. In our visual representation of digraphs, an arc $(i,j)$ is represented by an arrow starting from the initial vertex $i$ and ending at the terminal vertex $j$. The arc $(i,i)$ is called a self-loop at vertex ii. The outdegree of a vertex $v\in V$ is the number of arcs $(i,j)\in E$ such that $v=i$. Likewise, the indegree of a vertex $v\in V$ is the number of arcs $(i,j)\in E$ such that $v=j$. Let $G=(V,E)$ be a digraph and let $W\subseteq V$. The digraph induced by $W$ is the digraph $G|_{W}=(W,E|_{W})$ where $E|_{W}$ is the set of arcs $(i,j)$ such that $(i,j)\in E|_{W}$ if and only if $i,j\in W$.
Let $k\geq 0$ be an integer. A walk $\omega$ starting from vertex $i_{0}$ and ending at vertex $i_{k}$ in a digraph $G=(V,E)$ is a sequence of vertices $\langle i_{0},i_{1},\dots,i_{k}\rangle$ such that $(i_{t},i_{t+1})\in E$ for all $t\in\{0,1,\dots,k-1\}$. The length of the walk $\omega=\langle i_{0},i_{1},\dots,i_{k}\rangle$, denoted by $\lvert\omega\rvert$, is equal to $k$. We can also write the walk $\omega=\langle i_{0},i_{1},\dots,i_{k}\rangle$ as $i_{0}\to i_{1}\to\dots\to i_{k}$. Note that a walk may include a self-loop, or a vertex can be visited more than once. A walk that starts and ends at the same vertex is called a closed walk. For example, we consider any walk of length 0 to be a closed walk. A cycle is a closed walk of positive length with no repeated vertex. We consider a self-loop as a cycle of length 11. A directed graph with no cycle is called a directed acyclic graph (DAG).
Let $G=(V,E)$ be a digraph. We define a relation on $V$ where $i\sim j$ if and only if there is a walk in $G$ from $i$ to $j$, and there is also a walk from $j$ to ii. This relation is an equivalence relation on $V$. An equivalence class induced by this equivalence relation is called a strongly connected component (SCC) of $G$. If $G$ consists of only one SCC, then we call $G$ strongly connected. Equivalently, $G$ is strongly connected if and only if for all $i,j\in V$, there exists a walk in $G$ from $i$ to $j$. Let $W$ be an SCC of $G$. Clearly, the induced digraph $G|_{W}$ is strongly connected. We call $W$ a sink SCC if for all $(i,j)\in E$, $i\in W\implies j\in W$. In words, $W$ is a sink if there is no arc with its initial vertex in $W$ and its terminal vertex outside of $W$. For example, any vertex that has outdegree zero forms a sink SCC.
@@
Let AA be a row-stochastic matrix. The behavior of $\lim_{t\to\infty}A^{t}$ can be broadly categorized into two types:
@@
The following theorem is a combinatorial characterization of the existence of $\lim_{t\to\infty}A^{t}$ based on the underlying digraph of a given row-stochastic AA. Note that a row-stochastic matrix AA is condensely aperiodic if and only if for all sink SCC $W$ of $G[A]$, the induced digraph $G[A]\,|_{W}$ is aperiodic.
@@
Let AA be a row-stochastic matrix of positive size. Then $\lim_{t\to\infty}A^{t}=M$ where $M$ is row stochastic, if and only if AA is condensely aperiodic.
@@
Let AA be a row-substochastic matrix of positive size. If AA is condensely aperiodic then $\lim_{t\to\infty}A^{t}=M$ where $M$ is row substochastic.
@@
Suppose that AA is condensely aperiodic. Then for each $i$, $S_{i}=0$ or $S_{i}$ is aperiodic. Suppose $S_{i}$ is aperiodic and thus, primitive. Suppose also that $S_{i}$ is row stochastic, i.e., $\rho(S_{i})=1$. By the Perron–Frobenius theorem [48, Theorem 1.1], $S_{i}$ has 11 as a simple eigenvalue, and for all eigenvalues $\lambda\neq 1$ of $S_{i}$, we have $\lvert\lambda\rvert<1$. The eigenvalues of AA are precisely the eigenvalues of $U_{1},U_{2},\dots,U_{k},S_{1},S_{2},\dots,S_{m}$. Let $\lambda$ be an eigenvalue of AA. Obviously, $\lvert\lambda\rvert\leq 1$. Suppose that $\lvert\lambda\rvert=1$. Then at least one of $S_{1},S_{2},\dots,S_{m}$ has $\lambda$ as an eigenvalue. Suppose that for some $i$, $S_{i}$ has $\lambda$ as an eigenvalue. Since $\lvert\lambda\rvert=1$ then $\rho(S_{i})=1$. This implies that $S_{i}$ must be aperiodic and row stochastic. Hence, it is clear that $\lambda=1$. Furthermore, as an eigenvalue of AA, $\lambda=1$ is semisimple: its algebraic and geometric multiplicities are equal. By [31, Theorem 2.1], we conclude that there exists a matrix $M$ such that $\lim_{t\to\infty}A^{t}=M$. Moreover, observe that all entries of the vector $A^{t}\mathbf{1}$ is at most 11, which is denoted as $A^{t}\mathbf{1}\leq\mathbf{1}$. Taking limit yields $M\mathbf{1}\leq\mathbf{1}$ and thus, $M$ is also row substochastic. ∎
@@ 2 Time-invariant case and the spectral radius
In the time-invariant case, we have that $A_{t}=A$ and $\mathcal{E}_{t}=\mathcal{E}$ for all integers $t\geq 0$. We fix the notation AA and $\mathcal{E}$ for the rest of Section 2, unless stated otherwise. Next, Equation (2.1) becomes
@@
By Theorem 1.1, AA-$\mathcal{E}$ is zero-convergent if and only if $\rho(A-\mathcal{E})<1$. Even if $\rho(A-\mathcal{E})\geq 1$, it is possible that consensus can still be eventually reached, depending on $X_{0}-\bar{\sigma}$. An example is if $\lim_{t\to\infty}(A-\mathcal{E})^{t}=M$, where $M\neq 0$ is singular. Our main focus in this section is to derive various conditions for the entries of AA and $\mathcal{E}$ such that $\rho(A-\mathcal{E})<1$.
@@
We know that $\rho(A-\mathcal{E})\leq\lVert A$-$\mathcal{E}\rVert_{\infty}$. Suppose that $0<\mathcal{E}(i)<2A(i,i)$ for all $i$ in $\{1,2,\dots,n\}$. We claim that $\lVert A-\mathcal{E}\rVert_{\infty}<1$. Since AA is row stochastic, for each row $i\in\{1,2,\dots,n\}$, we have that
@@ 2.1 Anchor agents and the condensely-anchored property
Let AA be a row-stochastic matrix and let $\mathcal{E}$ be a diagonal nonnegative matrix of the same size $n>0$. Let $i\in\{1,2,\dots,n\}$. An agent $i$ is called an anchor of AA-$\mathcal{E}$ if $0<\mathcal{E}(i)<2A(i,i)$. Alternatively, the row/index/vertex $i$ is also called an anchor of AA-$\mathcal{E}$ if $0<\mathcal{E}(i)<2A(i,i)$. If $\mathcal{E}(i)=0$ then an agent $i$ is called defective. Consequently, an agent $i$ is a non-anchor if either $i$ is defective or $\mathcal{E}(i)\geq 2A(i,i)$.
@@
Let AA be a row-stochastic matrix of size $n>0$, and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$ such that $A-\mathcal{E}\geq 0$. Let $i\in\{1,2,\dots,n\}$. The following are equivalent:
@@
    The $i$-th row-sum of AA-$\mathcal{E}$ is strictly less than 11.
@@
Let $B$ be a row-substochastic matrix of size $n>0$. A row/index $i\in\{1,2,\dots,n\}$ is an anchor of $B$ if and only if the $i$-th row-sum of $B$ is strictly less than 11.
@@
Observe that there exists a unique pair of row-stochastic matrix AA and diagonal nonnegative matrix $\mathcal{E}$, both of size $n$, such that $B=A$-$\mathcal{E}$. The conclusion then immediately follows from $(1)\iff(3)$ in Proposition 2.4. ∎
By Corollary˜2.5, it is clear that any row-stochastic matrix AA does not have any anchor. Outside the context of this paper, Corollary˜2.5 can instead be used as the starting point to define anchors for row-substochastic matrices.
@@
Let AA be a row-stochastic matrix and let $\mathcal{E}$ be a diagonal nonnegative matrix of the same size. We call AA-$\mathcal{E}$ condensely anchored if every sink SCC of $G[A-\mathcal{E}]$ contains an anchor of AA-$\mathcal{E}$.
Note that Definition 2.6 above parallels that of condensely aperiodic. Clearly, if each row of AA-$\mathcal{E}$ is an anchor, then AA-$\mathcal{E}$ is condensely anchored. On the other hand, if AA-$\mathcal{E}$ does not have any anchor, then AA-$\mathcal{E}$ is not condensely anchored. For example, the row-stochastic matrix AA, if written as A0A-0, is not condensely anchored. Hence, this emphasizes that the definition of condensely anchored applies to AA-$\mathcal{E}$ instead of AA.
@@
Let AA be a row-stochastic matrix and let $\mathcal{E}$ be a diagonal nonnegative matrix of the same size $n>0$. Then AA-$\mathcal{E}$ is condensely anchored if and only if for all $i\in\{1,2,\dots,n\}$, there exists a walk in the digraph $G[A-\mathcal{E}]$ from vertex $i$ to some anchor vertex of AA-$\mathcal{E}$.
@@
Let $G=(V,E)$ be a digraph. Instead of involving anchors as defined in Definition˜2.3, we can simply fix any subset $U\subseteq V$. Hence, we can jointly reformulate both Definition 2.6 and Proposition 2.7 as: Every sink SCC of $G$ contains a vertex from $U$ if and only if for all $i\in V$, there exists a walk in $G$ from $i$ to some vertex in $U$. This statement is inherently graph-theoretic, as it can be stated independently of the matrices AA and $\mathcal{E}$.
@@ 2.2 The case where AA-$\mathcal{E}$ is nonnegative
Let $B$ be a row-substochastic matrix of size $n>0$. Let $i\in\{1,2,\dots,n\}$ and let $\widehat{W}_{i}(B)$ be the set of walks $\omega$ in $G[B]$ such that $\omega$ starts at $i$ and ends at some vertex $j$, where the $j$-th row-sum of $B$ is strictly less than 11. We define the index of contraction of $B$ as
@@
be a row-substochastic matrix. By Corollary˜2.5, note that only row 44 of $B$ is an anchor. There exists a unique pair of row-stochastic matrix
@@
Figure 7: The underlying digraph of both matrices AA and $B$ in Example˜2.10.
@@
Let AA be a row-stochastic matrix of size $n>0$, and let $\mathcal{E}$ be a diagonal nonnegative matrix of size $n$ such that $A-\mathcal{E}\geq 0$. Then, the following are equivalent:
@@
$(1)\iff(2)$ follows directly from Proposition˜2.7. By [4, Corollary 2.6] or [60, Theorem A.1.6], we obtain $(3)\iff(4)$. We will now show that $(2)\iff(3)$. First, we know that AA-$\mathcal{E}$ is row substochastic. From Definition˜2.9, we have that $\widehat{W}_{i}(A-\mathcal{E})$ is the set of walks $\omega$ in $G[A-\mathcal{E}]$ such that $\omega$ starts at $i$ and ends at some vertex $j$, where the $j$-th row-sum of AA-$\mathcal{E}$ is strictly less than 11. By Proposition 2.4, we also have that a vertex $j$ is an anchor of AA-$\mathcal{E}$ if and only if the $j$-th row-sum of AA-$\mathcal{E}$ is strictly less than 11. Hence, we obtain
@@
Let $B$ be a row-substochastic matrix of size $n>0$. We know that there exists a unique pair of row-stochastic matrix AA and diagonal nonnegative matrix $\mathcal{E}$, both of size $n$, such that $B=A$-$\mathcal{E}$. Hence, Theorem˜2.11 can also be stated in terms of row-substochastic matrix $B$ instead of AA-$\mathcal{E}$.
@@
In Example 2.13, we fix a row-stochastic matrix AA, and we consider two diagonal nonnegative matrices $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$. For each matrix $A-\mathcal{E}_{1}$ and $A-\mathcal{E}_{2}$, there are four defective agents but only a single anchor.
@@
The digraphs $G[A]$, $G[A-\mathcal{E}_{1}]$, and $G[A-\mathcal{E}_{2}]$ are all identical and condensely aperiodic, see Table˜1. Moreover, $\{1\}$ is the only sink SCC of each of these digraphs. Vertex 11 of $G[A-\mathcal{E}_{1}]$ is its sole anchor, while vertex 55 of $G[A-\mathcal{E}_{2}]$ is its sole anchor. In Table 1, we can see that $A-\mathcal{E}_{1}$ is condensely anchored, while $A-\mathcal{E}_{2}$ is not condensely anchored. By Theorem 2.11, we have $\lim_{t\to\infty}(A-\mathcal{E}_{1})^{t}=0$, while $A-\mathcal{E}_{2}$ is not zero-convergent. We find that